OpenStudy (anonymous):

A solenoid with 100 turns and a cross sectional area of 20cm^2 has a magnetic field of 0.60 T along its axis. If the field is confined within the solenoid and changes at a rate of 0.20 T/s, what is the magnitude of the induced EMF in the solenoid? I got e=4V but looks like that is wrong, can any one help pelase. Thanks!

OpenStudy (anonymous):

\[100 * 20/10^4 * 0.2\] = 0.04 V

OpenStudy (anonymous):

@zaphodplaysitsafe how did you arrive at that (which formula)??

OpenStudy (anonymous):

Induced EMF = rate of change of flux here flux = B * N * A where N is the number of turns so rate of change of flux here is = rate of change of B * N * A hence EMF = dB/dt * N * A

OpenStudy (anonymous):

@zaphodplaysitsafe thanks, looks like I forgot to square the conversion factor as I did 100*.2*.2. It's always the simple things.

OpenStudy (anonymous):

yeah thats true, it happens